# When Are Combinations of Humans and AI Useful?

### v1

@@
When Are Combinations of Humans and AI Useful?
@@
Inspired by the increasing use of AI to augment humans, researchers have studied human-AI systems involving different tasks, systems, and populations. Despite such a large body of work, we lack a broad conceptual understanding of when combinations of humans and AI are better than either alone. Here, we addressed this question by conducting a meta-analysis of over 100 recent experimental studies reporting over 300 effect sizes. First, we found that, on average, human-AI combinations performed significantly worse than the best of humans or AI alone. Second, we found performance losses in tasks that involved making decisions and significantly greater gains in tasks that involved creating content. Finally, when humans outperformed AI alone, we found performance gains in the combination, but when the AI outperformed humans alone we found losses. These findings highlight the heterogeneity of the effects of human-AI collaboration and point to promising avenues for improving human-AI systems.
@@ 1 Introduction
Extending the concepts of strong and weak synergy from human teams [larson2013search, almaatouq2021task], we focused on two types of synergy: (1) strong synergy, where the human-AI group performs better than both the human alone and the AI alone; and (2) weak synergy, where the human-AI group performs better than either the human or the AI alone but not both.
When evaluating human-AI systems, many studies focus on a specific kind of weak synergy where the baseline is the performance of the humans alone [bo2021toward, boskemper2022measuring, bondi2022role, schemmer2022meta]. This measure can serve important purposes in contexts for which full automation cannot happen for legal, ethical, safety, or other reasons. But when talking about the potential of human-AI systems, most people implicitly assume that the combined system should be better than either alone; otherwise, they would just use the best of the two [wilson2018collaborative]. Thus they are looking for strong synergy. In light of these considerations, a growing body of work emphasizes evaluating and searching for strong synergy in human-AI systems [bansal2021does, bansal2021most, bansal2019updates, wilder2020learning, rastogi2023taxonomy, mozannar2024effective].
To evaluate this synergy in human-AI systems, we analyzed 370 unique effect sizes from 106 different experiments published between January 2020 and June 2023 that included the performance of the human-only, AI-only, and human-AI systems. On average, we found evidence of weak synergy, meaning that the average human-AI systems performed better than the human alone. But we did not find strong synergy on average, meaning that the average human-AI systems performed worse than either the human alone or the AI alone. So, in practice, if we consider only the performance dimensions the researchers studied, it would have been better to use either a human alone or an AI system alone rather than the human-AI systems studied.
While this overall result may appear discouraging, we also identified specific factors that did or did not contribute to strong synergy in human-AI systems. On the one hand, for example, much of the recent research in human-AI collaboration has focused on using AI systems to help humans make decisions by providing not only suggested decisions but also confidence levels or explanations. But we found that neither of these factors significantly affected the performance of human-AI systems.
@@ 2 Results
Our initial literature search yielded 5126 papers, and, per the review process described in the Methods (see section 4.1), we identified 74 that met our inclusion criteria (see Figure S1). These papers reported the results of 106 unique experiments, and many of the experiments had multiple conditions, so we collected a total of 370 unique effect sizes measuring the impact of human-AI collaboration on task performance. Figure S2 highlights the descriptive statistics for the experiments in our analysis. We synthesize these data in a three-level meta-analytic model (see the Methods section 4.2). We also make the materials required to reproduce our results publicly accessible through an Open Science Framework (OSF) repository.
@@ 2.1 Overall Levels of Human-AI Synergy
In our primary analyses, we focus on strong human-AI synergy, so we compare the performance of the human-AI systems to a baseline of the human alone or the AI alone, whichever performed best. We found that the human-AI systems performed significantly worse overall than this baseline. The overall pooled effect was negative ($g=-0.23$, $p=0.0047$, 95% confidence interval (CI) $-0.39$ to $-0.07$) and considered small according to conventional interpretations [cohen2013statistical].
On the other hand, when we compared the performance of the human-AI systems to a different baseline – the humans alone – we found substantial evidence of weak human-AI synergy. The human-AI systems performed significantly better than humans alone, and this pooled effect size was positive ($g=0.64$, $p<0.0001$, 95% CI $0.53$ to $0.74$) and medium to large in size [cohen2013statistical]. Figure 1 displays a forest plot of these effect sizes. In other words, the human-AI systems we analyzed were, on average, better than humans alone but not better than both humans alone and AI alone. For effect sizes that correspond to other types of human-AI synergy, see Figure S3.
Figure 1: Forest plots of all effect sizes ($k=370$) included in the meta-analysis. The positions of the points on the $x$-axes represent the values of the effect sizes, and the bars indicate the 95% confidence interval for the effect sizes. The colors of the points and lines correspond to the values of the effect sizes, with negative effect sizes colored red (no human-AI synergy) and positive effect sizes colored green (human-AI synergy). The black dotted line corresponds to an effect size of Hedges’ $g=0$, which means that the human-AI group performed the same as the baseline. The circle at the bottom of the graph represents the meta-analytic average effect size and confidence interval.
@@ 2.2 Heterogeneity of Human-AI Synergy
We also found evidence for substantial heterogeneity of effect sizes in our estimations of both strong synergy ($I^{2}=97.7\%$) and weak synergy ($I^{2}=93.8\%$) (see table Table S4 for more details). Our moderator analysis identified characteristics of participants, tasks, and experiments that led to different levels of human-AI synergy, and it helps explain sources of this heterogeneity. Figure 2 provides a visualization of the results of meta-regressions with our moderators. Definitions of the subgroups for different moderator variables are included in Table S2, and more details of the regressions are included in Table S3.
Figure 2: Results from the three-level meta-regression models for the moderator variables. Here, $K$ is the number of included effect sizes for the moderator subgroup level, and the estimated effect size with the corresponding 95% confidence interval. The symbols in front of the moderator indicate if there is a statistically significant difference between the subgroups for strong synergy (*) and weak synergy ().
@@
First, we found that the type of task significantly moderated strong human-AI synergy ($p=0.0061$). Among decision tasks – those in which participants decided between a finite set of options – the pooled effect size for strong synergy was significantly negative ($g=-0.27$, 95%CI $-0.44$ to $-0.10$), which indicates performance losses from combining humans and AI. In contrast, among creation tasks – those in which participants created some sort of open response content – the pooled effect size for strong synergy was positive ($g=0.19$, 95%CI $-0.09$ to $0.48$), pointing to strong synergy between humans and AI. Even though the average performance gains for creation tasks were not significantly different from 0 (presumably because of the relatively small sample size of $n=34$), the difference between losses for decision tasks and gains for creation tasks was statistically significant. Relatedly, we found that the type of data involved in the task significantly moderated both strong and weak synergy in human-AI systems ($p<0.0001$).
@@
Second, we found that the performance of the human and AI relative to each other impacted both types of human-AI synergy ($p<0.0001$). As shown in Figure 2, when the human alone outperformed the AI alone, the combined human-AI system outperformed both alone with an average pooled effect size for strong synergy of $g=0.46$ (95%CI $0.28$ to $0.66$, a medium-sized effect [cohen2013statistical]). But when the AI alone outperformed the human alone, performance losses occurred in the combined system relative to the AI alone, with a negative effect size for strong synergy of $g=-0.54$ (95%CI $-0.71$ to $-0.37$, a medium -sized effect [cohen2013statistical]).
@@
The performance of the human and AI relative to each other also affected the degree of weak synergy in the human-AI systems. When the AI outperformed the human alone, greater performance gains tended to occur in the human-AI systems relative to the human alone ($p<0.0001$), and the pooled effect size for this type of weak synergy was positive and medium to large in magnitude ($g=0.74$, 95%CI $0.63$ to $0.85$) (see Figures S4 and S5 for a more detailed visualization of this result for decision tasks).
@@
Additionally, we found that the type of AI involved in the experiment moderated strong synergy in human-AI systems ($p=0.0264$), and the experimental design moderated weak synergy in human-AI systems ($p=0.0290$).
@@
The remaining moderators we investigated were not statistically significant for strong or weak human-AI synergy (explanation, confidence, participant type, division of labor).
@@ 3 Discussion
Performance Losses from Human-AI Collaboration: Regarding the first question, we found that, on average among recent experiments, human-AI systems did not exhibit strong synergy: the human-AI groups performed worse than either the human alone or the AI alone. This result complements the qualitative literature reviews on human-AI collaboration [lai2023towards, sperrle2021survey, maadi2021review], which highlight some of the surprising challenges that arise when integrating human and artificial intelligence. For example, people often rely too much on AI systems ("overreliance"), using its suggestions as strong guidelines without seeking and processing more information [skitka1999does, buccinca2021to, lai2020why]. Other times, however, humans rely too little on AI ("underreliance"), ignoring its suggestions because of adverse attitudes toward automation [zhang2020effect, buccinca2021to, vasconcelos2023explanations].
Interestingly, we found that, among this same set of experiments, weak synergy (better than humans) did exist in the human-AI systems: the human-AI groups performed better than the humans working alone. Thus, even though, on average, the human-AI combinations did not achieve strong synergy, they did, on average, augment human performance.
Moreover, the discrepancy between strong and weak synergy in human-AI groups mirrors that found in the literature on synergy in human-only groups. In both cases, substantial empirical evidence points to the existence of weak synergy, but there is relatively spare evidence of strong synergy in task performance [larson2013search, almaatouq2021task]. This result can occur, of course, because by definition the baseline for strong synergy is more stringent than that for weak synergy. It may also occur, however, because obtaining strong human-AI synergy requires different forms of human-AI interaction, or because the recent empirical studies were not appropriately designed to elicit strong synergy.
Moderating Effect of Task Type: With the large data set we collected, we also performed additional analyses of factors that influence the effectiveness of human-AI collaboration. We found that the type of task significantly moderated strong synergy in human-AI systems with decision tasks associated with performance losses and creation tasks associated with performance gains.
@@
With most of the decision tasks studied in our sample, however, both the human and the AI system make a complete decision, with the humans usually making the final choice. Our results suggest that with these decision tasks, better results might have been obtained if the experimenters had designed processes in which the AI systems did only the parts of the task for which they were clearly better than humans. Only 3 of the 100+ experiments in our analysis explore such processes with a pre-determined delegation of separate sub-tasks to humans and AI. With the 4 effect sizes from these 3 experiments, we found that, on average, strong human-AI synergy ($g=0.22,N.S.$) occurred (see Section S3.3 for a more detailed discussion of these experiments).
Moderating Effect of Relative Human/AI Performance: Interestingly, when the AI alone outperformed the human alone, substantial performance losses occurred in the human-AI systems ($g=-0.54$). When the human outperformed the AI alone, however, performance gains occurred in the human-AI systems ($g=0.46$). This finding shows that the human-AI performance cannot be explained with a simple average of the human alone and AI alone. In such a case, strong human-AI synergy could never exist [donahue2022human].
@@
Develop Generative AI for Creation Tasks. In our broad sample of recent experiments, the vast majority (about 85%) of the effect sizes were for decision-making tasks in which subjects chose among a predefined set of options. But in these cases we found that the average effect size for strong synergy was significantly negative. On the other hand, only about 10% of the effect sizes researchers studied were for creation tasks—those that involved open-ended responses. And in these cases we found that the average effect size for strong synergy was positive and significantly greater than that for decision tasks. This result suggests that studying human-AI synergy for creation tasks—many of which can be done with generative AI—could be an especially fruitful area for research.
@@
Develop More Robust Evaluation Criteria for Human-AI Systems. Many of the experiments in our analysis evaluate performance according to a single measure of overall accuracy, but this measure corresponds to different things depending on the situation, and it omits other important criteria for human-AI systems. For example, as one approaches the upper bound of performance, such as 100% accuracy, the improvements necessary to increase performance usually become more difficult for both humans and AI systems. In these cases, we may wish to consider a metric that applies a non-linear scaling to the overall classification accuracy and thus takes such considerations into account [campero2022test].
More importantly, there are many practical situations where good performance depends on multiple criteria. For instance, in many high-stakes settings such as radiology diagnoses and bail predictions, relatively rare errors may have extremely high financial or other costs. In these cases, even if AI can, on average, perform a task more accurately and less expensively than humans, it may still be desirable to include humans in the process if the humans are able to reduce the number of rare but very undesirable errors. One potential approach for situations like these is to create composite performance measures that incorporate the expected costs of various kinds of errors. The weak synergy measure described is also appropriate for these high-stakes settings.
@@ 4.1.3 Data Collection and Coding
We conducted the search in each of these databases in July 2023. To calculate our primary outcome of interest – the effect of combining human and artificial intelligence on task performance – we recorded the averages and standard deviations of the task performance of the human alone, the AI alone, and the human and AI working with each other, as well as the number of subjects in each of these conditions (see S2.2 for details).

### v2

@@
When Are Combinations of Humans and AI Useful? - A Systematic Review and Meta-Analysis
@@
Inspired by the increasing use of AI to augment humans, researchers have studied human-AI systems involving different tasks, systems, and populations. Despite such a large body of work, we lack a broad conceptual understanding of when combinations of humans and AI are better than either alone. Here, we addressed this question by conducting a pre-registered systematic review and meta-analysis of over 100 recent experimental studies reporting over 300 effect sizes. First, we found that, on average, human-AI combinations performed significantly worse than the best of humans or AI alone (Hedges’ $g=-0.23$, 95% confidence interval $-0.39$ to $-0.07$). Second, we found performance losses in tasks that involved making decisions and significantly greater gains in tasks that involved creating content. Finally, when humans outperformed AI alone, we found performance gains in the combination, but when the AI outperformed humans alone we found losses. These findings highlight the heterogeneity of the effects of human-AI collaboration and point to promising avenues for improving human-AI systems.
@@ 1 Introduction
We focused on two outcomes: (1) human-AI synergy, where the human-AI group performs better than both the human alone and the AI alone, which is analogous to strong synergy in human groups [larson2013search, almaatouq2021task]; and (2) human augmentation, where the human-AI group performs better than the human alone (see Section S1.1 in the SI for more details).
When evaluating human-AI systems, many studies focus on human augmentation [bo2021toward, boskemper2022measuring, bondi2022role, schemmer2022meta]. This measure can serve important purposes in contexts for which full automation cannot happen for legal, ethical, or safety reasons and in cases when AI does not align with human values. But when talking about the potential of human-AI systems, most people implicitly assume that the combined system should be better than either alone; otherwise, they would just use the best of the two [wilson2018collaborative]. Thus they are looking for human-AI synergy. In light of these considerations, a growing body of work emphasizes evaluating and searching for synergy in human-AI systems [bansal2021does, bansal2021most, bansal2019updates, wilder2020learning, rastogi2023taxonomy, mozannar2024effective].
To evaluate this synergy in human-AI systems, we analyzed 370 unique effect sizes from 106 different experiments published between January 2020 and June 2023 that included the performance of the human-only, AI-only, and human-AI systems. On average, we found evidence of human augmentation, meaning that the average human-AI systems performed better than the human alone. But we did not find human-AI synergy on average, meaning that the average human-AI systems performed worse than at least one of the human alone or the AI alone. So, in practice, if we consider only the performance dimensions the researchers studied, it would have been better to use either a human alone or an AI system alone rather than the human-AI systems studied.
While this overall result may appear discouraging, we also identified specific factors that did or did not contribute to synergy in human-AI systems. On the one hand, for example, much of the recent research in human-AI collaboration has focused on using AI systems to help humans make decisions by providing not only suggested decisions but also confidence levels or explanations. But we found that neither of these factors significantly affected the performance of human-AI systems.
@@ 2 Results
Our initial literature search yielded 5126 papers, and, per the review process described in the Methods (see section 4.1), we identified 74 that met our inclusion criteria (see Figure S1). These papers reported the results of 106 unique experiments, and many of the experiments had multiple conditions, so we collected a total of 370 unique effect sizes measuring the impact of human-AI collaboration on task performance. Figure S2 highlights the descriptive statistics for the effect sizes in our analysis. We synthesize these data in a three-level meta-analytic model (see the Methods section 4.2). We also make the materials required to reproduce our results publicly accessible through an Open Science Framework (OSF) repository.
@@ 2.1 Overall Levels of Human-AI Synergy
In our primary analyses, we focus on human-AI synergy, so we compare the performance of the human-AI systems to a baseline of the human alone or the AI alone, whichever performed best. We found that the human-AI systems performed significantly worse overall than this baseline. The overall pooled effect was negative ($g=-0.23$, $t(92)=-2.89$, two-tailed $p=0.005$, 95% confidence interval (CI) $-0.39$ to $-0.07$) and considered small according to conventional interpretations [cohen2013statistical].
On the other hand, when we compared the performance of the human-AI systems to a different baseline – the humans alone – we found substantial evidence of human augmentation. The human-AI systems performed significantly better than humans alone, and this pooled effect size was positive ($g=0.64$, $t(98)=11.87$, two-tailed $p=0.000$, 95% CI $0.53$ to $0.74$) and medium to large in size [cohen2013statistical]. Figure 1 displays a forest plot of these effect sizes. In other words, the human-AI systems we analyzed were, on average, better than humans alone but not better than both humans alone and AI alone. For effect sizes that correspond to other potential outcomes of interest, see Table S3 and Figure S3.
@@ 2.2 Heterogeneity of Human-AI Synergy
We also found evidence for substantial heterogeneity of effect sizes in our estimations of human-AI synergy ($I^{2}=97.7\%$) and human augmentation ($I^{2}=93.8\%$) (see table Table S5 and S6 for more details). Our moderator analysis identified characteristics of participants, tasks, and experiments that led to different levels of human-AI synergy and human augmentation, and it helps explain sources of this heterogeneity. Figure 2 provides a visualization of the results of meta-regressions with our moderators. Definitions of the subgroups for different moderator variables are included in Table S2, and more details of the regressions for other potential outcomes of interest are included in Table S7.
@@
First, we found that the type of task significantly moderated human-AI synergy ($F(1,104)=7.84$, two-tailed $p=0.006$). Among decision tasks—those in which participants decided between a finite set of options—the pooled effect size for human-AI synergy was significantly negative ($g=-0.27$, $t(104)=-3.20$, two-tailed $p=0.002$, 95%CI $-0.44$ to $-0.10$), which indicates performance losses from combining humans and AI. In contrast, among creation tasks—those in which participants created some sort of open response content—the pooled effect size for human-AI synergy was positive ($g=0.19$, $t(104)=1.35$, two-tailed $p=0.180$, 95%CI $-0.09$ to $0.48$), pointing to synergy between humans and AI. Even though the average performance gains for creation tasks were not significantly different from 0 (presumably because of the relatively small sample size of $n=34$), the difference between losses for decision tasks and gains for creation tasks was statistically significant. Relatedly, we found that the type of data involved in the task significantly moderated both human-AI synergy ($F(4,101)=15.24$, two-tailed $p=0.000)$ and human augmentation ($F(4,101)=6.52$, two-tailed $p=0.000)$).
@@
Second, we found that the performance of the human and AI relative to each other impacted both human-AI synergy ($F(1,104)=81.79$, two-tailed $p=0.000$) and human augmentation ($F(1,104)=24.35$, two-tailed $p=0.000$). As shown in Figure 2, when the human alone outperformed the AI alone, the combined human-AI system outperformed both alone with an average pooled effect size for human-AI synergy of $g=0.46$ ($t(104)=5.06$, two-tailed $p=0.000$, 95%CI $0.28$ to $0.66$), a medium-sized effect [cohen2013statistical]). But when the AI alone outperformed the human alone, performance losses occurred in the combined system relative to the AI alone, with a negative effect size for human-AI synergy of $g=-0.54$ ($t(104)=-6.20$, two-tailed $p=0.000$, 95%CI $-0.71$ to $-0.37$), a medium -sized effect [cohen2013statistical]). Given the importance of this moderator, we fit separate meta-analytic models on the subset of results where (1) the AI performs better alone and (2) the human performs better alone, and we report the results for human-AI synergy and human augmentation in Table S4.
@@
The performance of the human and AI relative to each other also affected the degree of human augmentation in the human-AI systems ($F(1,104)=24.35$, two-tailed $p=0.000$). When the AI outperformed the human alone, greater performance gains tended to occur in the human-AI systems relative to the human alone, and the pooled effect size for human augmentation was positive and medium to large in magnitude ($g=0.74$, $t(104)=13.50$, two-tailed $p=0.000$, 95%CI $0.63$ to $0.85$) (see Figures S4-S7 for a more detailed visualization of this result for decision tasks).
@@
Additionally, we found that the type of AI involved in the experiment ($F(2,103)=3.77$, two-tailed $p=0.026$) and the year of publication ($F(3,102)=3.65$, two-tailed $p=0.015$) moderated human-AI synergy, and the experimental design moderated human augmentation ($F(1,104)=4.90$, two-tailed $p=0.029$). See Figure S8 for a more detailed visualization of the effect sizes by year of publication.
@@
The remaining moderators we investigated were not statistically significant for human-AI synergy or human augmentation (explanation, confidence, participant type, division of labor).
@@ 3 Discussion
Performance Losses from Human-AI Collaboration: Regarding the first question, we found that, on average among recent experiments, human-AI systems did not exhibit synergy: the human-AI groups performed worse than either the human alone or the AI alone. This result complements the qualitative literature reviews on human-AI collaboration [lai2023towards, sperrle2021survey, maadi2021review], which highlight some of the surprising challenges that arise when integrating human and artificial intelligence. For example, people often rely too much on AI systems ("overreliance"), using its suggestions as strong guidelines without seeking and processing more information [skitka1999does, buccinca2021to, lai2020why]. Other times, however, humans rely too little on AI ("underreliance"), ignoring its suggestions because of adverse attitudes toward automation [zhang2020effect, buccinca2021to, vasconcelos2023explanations].
Interestingly, we found that, among this same set of experiments, human augmentation did exist in the human-AI systems: the human-AI groups performed better than the humans working alone. Thus, even though, on average, the human-AI combinations did not achieve synergy, the AI system did, on average, help humans perform better. This result can occur, of course, because by definition the baseline for human-AI synergy is more stringent than that for human augmentation. It may also occur, however, because obtaining human-AI synergy requires different forms of human-AI interaction, or because the recent empirical studies were not appropriately designed to elicit human-AI synergy.
Moderating Effect of Task Type: With the large data set we collected, we also performed additional analyses of factors that influence the effectiveness of human-AI collaboration. We found that the type of task significantly moderated synergy in human-AI systems with decision tasks associated with performance losses and creation tasks associated with performance gains.
@@
With most of the decision tasks studied in our sample, however, both the human and the AI system make a complete decision, with the humans usually making the final choice. Our results suggest that with these decision tasks, better results might have been obtained if the experimenters had designed processes in which the AI systems did only the parts of the task for which they were clearly better than humans. Only 3 of the 100+ experiments in our analysis explore such processes with a pre-determined delegation of separate sub-tasks to humans and AI. With the 4 effect sizes from these 3 experiments, we found that, on average, human-AI synergy ($g=0.22$, $t(104)=0.69$, two-tailed $p=0.494$, 95% CI $-0.42$ to $0.87$) occurred, but the result was not statistically significant (see Section S2.6 for a more detailed discussion of these experiments).
Moderating Effect of Relative Human/AI Performance: Interestingly, when the AI alone outperformed the human alone, substantial performance losses occurred in the human-AI systems. When the human outperformed the AI alone, however, performance gains occurred in the human-AI systems. This finding shows that the human-AI performance cannot be explained with a simple average of the human alone and AI alone. In such a case, human-AI synergy could never exist [donahue2022human].
@@
Develop Generative AI for Creation Tasks. In our broad sample of recent experiments, the vast majority (about 85%) of the effect sizes were for decision-making tasks in which subjects chose among a predefined set of options. But in these cases we found that the average effect size for human-AI synergy was significantly negative. On the other hand, only about 10% of the effect sizes researchers studied were for creation tasks—those that involved open-ended responses. And in these cases we found that the average effect size for human-AI synergy was positive and significantly greater than that for decision tasks. This result suggests that studying human-AI synergy for creation tasks—many of which can be done with generative AI—could be an especially fruitful area for research.
@@
Develop More Robust Evaluation Metrics for Human-AI Systems. Many of the experiments in our analysis evaluate performance according to a single measure of overall accuracy, but this measure corresponds to different things depending on the situation, and it omits other important criteria for human-AI systems. For example, as one approaches the upper bound of performance, such as 100% accuracy, the improvements necessary to increase performance usually become more difficult for both humans and AI systems. In these cases, we may wish to consider a metric that applies a non-linear scaling to the overall classification accuracy and thus takes such considerations into account [campero2022test].
More importantly, there are many practical situations where good performance depends on multiple criteria. For instance, in many high-stakes settings such as radiology diagnoses and bail predictions, relatively rare errors may have extremely high financial or other costs. In these cases, even if AI can, on average, perform a task more accurately and less expensively than humans, it may still be desirable to include humans in the process if the humans are able to reduce the number of rare but very undesirable errors. One potential approach for situations like these is to create composite performance measures that incorporate the expected costs of various kinds of errors. The human augmentation measure described is also appropriate for these high-stakes settings.
@@ 4.1.3 Data Collection and Coding
We conducted the search in each of these databases in July 2023. To calculate our primary outcome of interest – the effect of combining human and artificial intelligence on task performance – we recorded the averages and standard deviations of the task performance of the human alone, the AI alone, and the human and AI working with each other, as well as the number of subjects in each of these conditions (see S1.3 for details).